# Region-Level Context-Aware Multimodal Understanding

### v1

@@ Abstract
Despite significant progress, existing research on Multimodal Large Language Models (MLLMs) mainly focuses on general visual understanding, overlooking the ability to integrate textual context associated with objects for a more context-aware multimodal understanding — an ability we refer to as Region-level Context-aware Multimodal Understanding (RCMU). To address this limitation, we first formulate the RCMU task, which requires models to respond to user instructions by integrating both image content and textual information of regions or objects. To equip MLLMs with RCMU capabilities, we propose Region-level Context-aware Visual Instruction Tuning (RCVIT), which incorporates object information into the model input and enables the model to utilize bounding box coordinates to effectively associate objects’ visual content with their textual information. To address the lack of datasets, we introduce the RCMU dataset, a large-scale visual instruction tuning dataset that covers multiple RCMU tasks. We also propose RC&P-Bench, a comprehensive benchmark that can evaluate the performance of MLLMs in RCMU and multimodal personalized understanding tasks. Additionally, we propose a reference-free evaluation metric to perform a comprehensive and fine-grained evaluation of the region-level context-aware image descriptions. By performing RCVIT on Qwen2-VL models with the RCMU dataset, we developed RC-Qwen2-VL models. Experimental results indicate that RC-Qwen2-VL models not only achieve outstanding performance on multiple RCMU tasks but also demonstrate successful applications in multimodal RAG and personalized conversation. Our data, model and benchmark are available at https://github.com/hongliang-wei/RC-MLLM
@@
Article submission, IEEE, IEEEtran, journal, , paper, template, typesetting.
@@ IV RC&P-Bench
RC&P-Bench contains 3269 multiple-choice questions, covering 259 entities.Each question image contains at least two personalized entities, and all personalized entities within a single image belong to the same category. The average length of personalized information for each entity is 128.1 words.
@@
C2V V2C Overall C2V V2C Overall C2V V2C Overall
GPT-4o Proprietary 38.18 45.15 46.75 62.04 66.27 65.18 67.79 78.51 71.29
Gemini Pro 2.5 64.06 64.31 68.56 63.09 59.58 66.23 77.18 77.39 79.23
InternVL2-2B Open-source 22.30 24.64 28.10 26.99 40.12 36.35 35.12 35.16 39.26
InternVL2-4B 20.66 19.68 27.09 53.84 48.38 52.45 47.35 34.41 45.06
InternVL2-8B 26.10 32.61 31.81 57.05 50.04 51.66 75.69 72.35 65.61
MiniCPM V2.6 (8B) 22.60 26.00 30.67 40.87 45.15 46.01 57.05 52.97 53.80
RAP-LLaVA 13B 4.77 5.11 4.97 6.26 6.46 6.13 7.31 7.21 7.27
Qwen2-VL 2B 24.01 31.03 25.18 14.84 32.91 22.18 41.61 53.34 39.66
RC-Qwen2-VL 2B (Ours) 29.38 +5.37 34.49 +3.46 30.43 +5.25 54.74 +39.90 46.13 +13.22 45.61 +23.43 62.12 +20.51 62.89 +9.55 52.76 +13.10
Qwen2-VL 7B 25.35 30.73 31.26 32.29 44.55 39.75 59.88 62.58 56.90
RC-Qwen2-VL 7B (Ours) 43.70 +18.35 43.65 +12.92 42.45 +11.19 70.54 +38.25 53.94 +9.39 57.70 +17.95 93.81 +33.93 88.58 +26.00 78.83 +21.93
@@ V-C Experimental Results on Personalized Multimodal Understanding
Results and Analysis. The results on RC&P-Bench are presented in Table III, showing that the RC-Qwen2-VL models achieve significant improvements over the baseline models. Specifically, RC-Qwen2-VL 2B shows a remarkable +23.43 boost in the Open-World Overall score and +13.10 in the Oracle setting compared to its Qwen2-VL 2B counterpart. The gains are even more pronounced for the larger model. RC-Qwen2-VL 7B surpasses its baseline by +17.95 in the Open-World Overall score and an impressive +21.93 in the Oracle setting. The most substantial individual improvements are seen in the Context-to-Visual (C2V) tasks, with RC-Qwen2-VL 7B gaining +38.25 in the Open-World C2V and +33.93 in the Oracle C2V tasks. Furthermore, the RC-Qwen2-VL 7B model has achieved state-of-the-art (SOTA) performance among open-source models and is comparable to or even surpasses top-tier proprietary models. For instance, in the Oracle setting, RC-Qwen2-VL 7B’s Overall score of 78.83% significantly outperforms other open-source models like InternVL2-8B (65.61%) and MiniCPM V2.6 (53.80%). Notably, this score also surpasses the proprietary GPT-4o (71.29%) and is highly competitive with Gemini Pro 2.5 (79.23%). This indicates that the capabilities of the RC-Qwen2-VL models, developed through region-level context-aware visual instruction tuning, can be effectively transferred to multimodal personalized understanding.

### v2

@@ Abstract
Despite significant progress, existing research on Multimodal Large Language Models (MLLMs) mainly focuses on general visual understanding, overlooking the ability to integrate textual context associated with objects for a more context-aware multimodal understanding – an ability we refer to as Region-level Context-aware Multimodal Understanding (RCMU). To address this limitation, we first formulate the RCMU task, which requires models to respond to user instructions by integrating both image content and textual information of regions or objects. To equip MLLMs with RCMU capabilities, we propose Region-level Context-aware Visual Instruction Tuning (RCVIT), which incorporates object information into the model input and enables the model to utilize bounding box coordinates to effectively associate objects’ visual content with their textual information. To address the lack of datasets, we introduce the RCMU dataset, a large-scale visual instruction tuning dataset that covers multiple RCMU tasks. We also propose RC&P-Bench, a comprehensive benchmark that can evaluate the performance of MLLMs in RCMU and multimodal personalized understanding tasks. Additionally, we propose a reference-free evaluation metric to perform a comprehensive and fine-grained evaluation of the region-level context-aware image descriptions. By performing RCVIT on Qwen2-VL models with the RCMU dataset, we developed RC-Qwen2-VL models. Experimental results indicate that RC-Qwen2-VL models not only achieve outstanding performance on multiple RCMU tasks but also demonstrate successful applications in multimodal RAG and personalized conversation. Our data, model and benchmark are available at https://github.com/hongliang-wei/RC-MLLM
@@
Multimodal Large Language Model, Region-Level Context-Aware Multimodal Understanding, Personalized Multimodal Understanding, Dataset, Benchmark
@@ IV RC&P-Bench
RC&P-Bench contains 3260 multiple-choice questions, covering 259 entities.Each question image contains at least two personalized entities, and all personalized entities within a single image belong to the same category. The average length of personalized information for each entity is 128.1 words.
@@
V2C C2V Overall V2C C2V Overall V2C C2V Overall
GPT-4o Proprietary 48.79 44.42 46.75 64.96 65.44 65.18 71.75 70.76 71.29
Gemini Pro 2.5 68.99 68.07 68.56 65.36 67.21 66.23 78.77 79.76 79.23
InternVL2-2B Open-source 30.72 25.10 28.10 42.58 29.24 36.35 39.99 38.44 39.26
InternVL2-4B 29.80 23.98 27.09 51.78 53.22 52.45 41.08 49.61 45.06
InternVL2-8B 35.39 27.73 31.81 48.73 54.99 51.66 60.47 71.48 65.61
MiniCPM V2.6 (8B) 34.81 25.95 30.67 49.48 42.05 46.01 51.32 56.64 53.80
RAP-LLaVA 13B 5.12 4.80 4.97 6.16 6.11 6.13 7.25 7.29 7.27
Qwen2-VL 2B 28.25 21.68 25.18 29.69 13.60 22.18 41.94 37.06 39.66
RC-Qwen2-VL 2B (Ours) 32.28 +4.03 28.32 +6.64 30.43 +5.25 41.20 +11.51 50.66 +37.06 45.61 +23.43 49.71 +7.77 56.24 +19.18 52.76 +13.10
Qwen2-VL 7B 34.06 28.06 31.26 44.65 34.17 39.75 54.95 59.13 56.90
RC-Qwen2-VL 7B (Ours) 41.83 +7.77 43.17 +15.11 42.45 +11.19 49.71 +5.06 66.82 +32.65 57.70 +17.95 71.29 +16.34 87.45 +28.32 78.83 +21.93
@@ V-C Experimental Results on Personalized Multimodal Understanding
Results and Analysis. The results on RC&P-Bench are presented in Table III, showing that the RC-Qwen2-VL models achieve significant improvements over the baseline models. Specifically, RC-Qwen2-VL 2B shows a remarkable 23.43 boost in the Open-World Overall score and 13.10 in the Oracle setting compared to its Qwen2-VL 2B baseline. The gains are even more pronounced for the larger model. RC-Qwen2-VL 7B surpasses its baseline by 17.95 in the Open-World Overall score and an impressive 21.93 in the Oracle setting. The most substantial individual improvements are seen in the Context-to-Visual (C2V) tasks, with RC-Qwen2-VL 7B gaining 32.65 in the Open-World C2V and 28.32 in the Oracle C2V tasks. Furthermore, the RC-Qwen2-VL 7B model has achieved state-of-the-art (SOTA) performance among open-source models and is comparable to or even surpasses top-tier proprietary models. For instance, in the Oracle setting, RC-Qwen2-VL 7B’s Overall score of 78.83% significantly outperforms other open-source models like InternVL2-8B (65.61%) and MiniCPM V2.6 (53.80%). Notably, this score also surpasses the proprietary GPT-4o (71.29%) and is highly competitive with Gemini Pro 2.5 (79.23%). This indicates that the capabilities of the RC-Qwen2-VL models, developed through region-level context-aware visual instruction tuning, can be effectively transferred to multimodal personalized understanding.